# Constraining the position of the CEP through the Speed of Sound in the LSMq

### 08 Feb 2024 v1

@@ Abstract
Examples of the shape of the phase diagram that we can draw from our perspective are depicted in Fig. 3 for values corresponding to $T_{c}^{0}$ for $N_{f}=2$ and $N_{f}=2+1$ flavors in LQCD simulations. We select the speed of sound $c_{s/\rho}=0.16$ for this exercise. We notice that the first order transition is achieved for small values of temperature, $T_{E}\simeq 80{\rm MeV}$, and values of the baryon chemical potential $\mu_{B}^{E}\simeq 800{\rm MeV}$ for $N_{f}=2$, but for $N_{f}=2+1$ the CEP cannot be located, as the high-temperature assumption is no longer reliable. Additionally, in Fig. 4 we sketch the phase diagram and the value of $c_{s/\rho}^{2}$ for all permitted values of the approximation. For $N_{f}=2$, the CEP can be located and a sharp fall-off of $c_{s/\rho}^{2}$ is noticeable. This is not the case of $N_{f}=2+1$, where no evidence of first order transition can be found.
Figure 4: effective phase diagram obtained with the parameters $a$, $\lambda$ and $g$ that ensure the value of $c_{s/\rho}^{2}=0.16$ at $T_{c}^{0}=166{\rm MeV}$ and $T_{c}^{0}=158{\rm MeV}$ with the value of $c_{s/\rho}^{2}$ for valid $T$ and $\mu_{B}$.
@@
In Fig. 5 we depict the behavior of $c_{s/\rho}^{2}$ for different critical values $\mu_{B}^{c}$ and the corresponding critical temperatures $T_{c}$. Again, different curves are for the parameters of $N_{f}=2$ and $N_{f}=2+1$ in LQCD. In accordance with the position of the CEP we distinguish $c_{s/\rho}^{2}$ in the continuous and discontinuous phase transitions only for the case of light quarks. The sharp drop-off of $c_{s/\rho}^{2}$ near the CEP can in fact be used as a signal to detect its position. For $N_{f}=2+1$, expectedly, the high-$T$ approximation breaks as well and thus, no evidence of a CEP is observed.
In this letter we re-examined the LSMq in connection with the physics of the QCD phase diagram and the existence of the CEP. We have put forward a strategy to fix the couplings $g$ and $\lambda$ as well as the parameter $a$ in the Lagrangian (1) from three requirements: (i) the condition of continuous phase transitions at high temperature, (5), the curvature of the critical line in the phase diagram according to LQCD, (6), and demanding that the speed of sound $c_{s/\rho}^{2}$ corresponds to the measured value. We consider that such a strategy allows to further constraint the free parameter space and discard some approximations in the need for more educated analysis of the phase diagram through effective models of this kind. As compared with a previous analysis by some of us regarding the collision energy dependence in the position of the CEP [30], we have narrowed down the range of values of $\lambda$ to a single point. We stress that the attainable values of $c_{s/\rho}$ in that work are significantly separated from the measured value. In this regard, our demand that the parameter choice renders the measured value of the speed of sound is in fact a tight constraint that any treatment of the effective potential of the LSMq should fulfill.

### 05 Mar 2024 v2

@@ Abstract
Examples of the shape of the phase diagram that we can draw from our perspective are depicted in Fig. 3 for values corresponding to $T_{c}^{0}$ for $N_{f}=2$ and $N_{f}=2+1$ flavors in LQCD simulations. We select the speed of sound $c_{s/\rho}=0.16$ for this exercise. We notice that the first order transition is achieved for small values of temperature, $T_{E}\simeq 80{\rm MeV}$, and values of the baryon chemical potential $\mu_{B}^{E}\simeq 800{\rm MeV}$ for $N_{f}=2$, but for $N_{f}=2+1$ the CEP cannot be located, as the high-temperature assumption is no longer reliable.
@@
In Fig. 4 we depict the behavior of $c_{s/\rho}^{2}$ for different critical values $\mu_{B}^{c}$ and the corresponding critical temperatures $T_{c}$. Again, different curves are for the parameters of $N_{f}=2$ and $N_{f}=2+1$ in LQCD. For $N_{f}=2$, the CEP can be located and a sharp fall-off of $c_{s/\rho}^{2}$ is noticeable which can be used as a signal to dectect its position. This is not the case of $N_{f}=2+1$, where no evidence of first order transition can be found because the hight-$T$ approximations breaks before reaching the CEP.
In this letter we re-examined the LSMq in connection with the physics of the QCD phase diagram and the existence of the CEP. We have put forward a strategy to fix the couplings $g$ and $\lambda$ as well as the parameter $a$ in the Lagrangian (1) from three requirements: (i) the condition of continuous phase transitions at high temperature, (5), (ii) the curvature of the critical line in the phase diagram according to LQCD, (6), and (iii) demanding that the speed of sound $c_{s/\rho}^{2}$ corresponds to the measured value. We consider that such a strategy allows to further constrain the free parameter space and discard some approximations in the need for more educated analysis of the phase diagram through effective models of this kind. As compared with a previous analysis by some of us regarding the collision energy dependence in the position of the CEP [30], we have narrowed down the range of values of $\lambda$ to a single point. We stress that the attainable values of $c_{s/\rho}$ in that work are significantly separated from the measured value. In this regard, our demand that the parameter choice renders the measured value of the speed of sound is in fact a tight constraint that any treatment of the effective potential of the LSMq should fulfill.